How much 90 percent vinegar do you have to add to a gallon of 5 percent vinegar to get 20 percent vinegar?
Answer by stanbon(75887)

How much 90 percent vinegar do you have to add to a gallon of 5 percent vinegar to get 20 percent vinegar?
--------
Equation:
vinegar + vinegar = vinegar
0.90x + 0.05*1 = 0.20(x+1)
-----
90x + 5 = 20x + 20
----
70x = 15
x = 15/70 = 0.21 gallons (amt. 90% solution needed)
------------------
Cheers,
Stan H.
==================